OpenStudy (anonymous):

Find the common solution of the equations - \[2\sin^{2}x+\sin^{2}2x=2 \]and\[\sin2x+\cos2x=tanx \]

OpenStudy (anonymous):

play around first one with\[\sin^2 2x=4\sin^2 x\cos^2x\]and replace \(\cos^2x\) with \(1-\sin^2 x\)

OpenStudy (anonymous):

For 1st eq.\[x=n \pi \pm \frac{\pi}{4}\] or\[x=n \pi \pm \frac{\pi}{2}\]

OpenStudy (anonymous):

For 2nd eq. \[x=\frac{n \pi}{2}-(-1)^{n}\frac{\pi}{4}\] or\[x={n \pi}{}\pm\frac{\pi}{4}\]

OpenStudy (anonymous):

@mukushla Now what to do??

OpenStudy (experimentx):

if you do that ... the first equation is quadratic in sin^2 x

OpenStudy (anonymous):

after solving both quadratics i got the above sol.

OpenStudy (anonymous):

ans. common sol.- \[x=(2n+1)\frac{\pi}{4}\]

OpenStudy (anonymous):

????

OpenStudy (anonymous):

i dont understand that solution

OpenStudy (experimentx):

how did you solve the second equation?

OpenStudy (anonymous):

\[2sinxcosx+1-2\sin^{2}x=\frac{sinx}{cosx}\]

OpenStudy (anonymous):

\[2sinxcos^{2}x+cosx-2\sin^{2}xcosx-{sinx}=0\]

OpenStudy (anonymous):

\[(2sinxcos+1)(cosx-sinx)=0\]

OpenStudy (experimentx):

pi/4 seems to be one solution ... what happened to the wolf http://www.wolframalpha.com/input/?i=solve+sin%282x%29%2B+cos%282x%29+%3D+tan%28x%29

OpenStudy (anonymous):

Can't we use tan formula there ??

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\frac{2tanx}{1 + \tan^2x} + \frac{1 - \tan^2x}{1 + \tan^2x} = tanx\]

OpenStudy (anonymous):

Or may be I am wrong too or may be I am going long by using that method..

OpenStudy (experimentx):

graphically the solution seems correct http://www3.wolframalpha.com/Calculate/MSP/MSP17141a3343heb1b3b16d00005b4iicdfaeaf7hhd?MSPStoreType=image/gif&s=43&w=425&h=188&cdf=Coordinates&cdf=Tooltips man what happen to wolf ... why is it giving crazy result

OpenStudy (experimentx):

\[ \sqrt 2 \sin\left( 2x + {\pi \over 4}\right) = \tan x\] the solution seems correct. maybe i'm using WA too much for free.

OpenStudy (experimentx):

Damn ... wolf was right ... sqrt(2) arctan(1 -sqrt(2)) = pi/4 (didn't know this identity) http://www.wolframalpha.com/input/?i=2+arctan%281+-+sqrt%282%29%29